A $9000 deposit in an account with an APR of 6.1% compounded continuously for 5 years.
Answer by stanbon(75887)

A $9000 deposit in an account with an APR of 6.1% compounded continuously for 5 years.
--------------
A(t) = P*e^(rt)
----
A(5) = 9000*e^(0.061*5) = $12209.63
------------
Cheers,
Stan H.
--------------